A teacher sits on a rotating stool to demonstrate the conservation of angular


momentum. With her arms outstretched, she rotates at 2.25 rev/s. When the


teacher’s arms are tucked in close to her chest, she rotates at 4.50 rev/s. what


is the ratio of her moment of inertia with outstretched arms to the moment of


inertia with arms tucked in close to her chest?

1
Expert's answer
2021-11-08T17:10:10-0500

angular momentum:

"L=\omega I"

where "\omega" is angular velocity, I is moment of inertia


"\omega =2\pi n",

where n is revolutions per second


conservation of angular momentum:

"\omega_1 I_1=\omega_2 I_2"


the ratio of her moment of inertia with outstretched arms to the moment of

inertia with arms tucked in close to her chest:


"\frac{I_1}{I_2}=\frac{\omega_2}{\omega_1}=\frac{n_2}{n_1}=2.25/4.5=1/2"